OpenStudy (anonymous):

write this equation in standard form using integers. y+3=4(x-1)

jigglypuff314 (jigglypuff314):

standard form of linear equation: ax + by = c (as in you want to get all the stuff with variables to one side) so for y+3=4(x-1) first distribute then you can subtract y from both sides then add 4 to both sides